OpenStudy (anonymous):

Will Medal & Fan If the mean of this data is 95, find the missing data. 103,99,90,86,105,96,__

OpenStudy (anonymous):

@midnighti

OpenStudy (solomonzelman):

The mean is the sum of values divided by the number of values.

OpenStudy (solomonzelman):

You can call the missing value - "x"

OpenStudy (solomonzelman):

And therefore you are having, \(\color{#000000 }{ \displaystyle \frac{103+9+9+86+105+96+x }{7} =95 }\) I am dividing by 7, because there are 7 values. And I set it equal to 95, because I know that the mean (or the "average") is 95.